# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning (RL) is a powerful approach for data-driven decision-making and control. Compared to model-free methods, offline model-based reinforcement learning (MBRL) explicitly learns world models from a static dataset and uses them as surrogate simulators, improving the data efficiency and enabling the learned policy to potentially generalize beyond the dataset support. However, there could be various MDPs that behave identically on the offline dataset and so dealing with the uncertainty about the true MDP can be challenging. In this paper, we propose modeling offline MBRL as a Bayes Adaptive Markov Decision Process (BAMDP), which is a principled framework for addressing model uncertainty. We further introduce a novel Bayes Adaptive Monte-Carlo planning algorithm capable of solving BAMDPs in continuous state and action spaces with stochastic transitions. This planning process is based on Monte Carlo Tree Search and can be integrated into offline MBRL as a policy improvement operator in policy iteration. Our “RL + Search" framework follows in the footsteps of superhuman AIs like AlphaZero, improving on current offline MBRL methods by incorporating more computation input. The proposed algorithm significantly outperforms state-of-the-art model-based and model-free offline RL methods on twelve D4RL MuJoCo benchmark tasks and three target tracking tasks in a challenging, stochastic tokamak control simulator.
@@ 1 Introduction
The success of reinforcement learning (RL) typically relies on large amounts of interactions with the environment. However, in real-world scenarios, such interactions can be unsafe or costly. As an alternative, offline RL (Levine et al. (2020)) leverages offline datasets of transitions, collected by a behavior policy, to train a policy that can transfer to an online task. To avoid overestimation of the value function for some (out-of-sample) states in the environment, which can mislead policy learning, model-free offline RL methods (Kumar et al. (2020); Wu et al. (2019)) often constrain the learned policy to remain close to the behavior policy or within the support of the offline dataset. However, collecting transitions that comprehensively cover possible task scenarios, or acquiring a large volume of demonstrations from a high-quality behavior policy, can be expensive. This challenge has led to the development of offline model-based reinforcement learning (MBRL) approaches, such as (Lu et al. (2022); Guo et al. (2022)). These methods train dynamics models from offline data and optimize policies using imaginary rollouts generated by the models. Notably, the dynamics modeling is independent of the behavior policy, making it possible to achieve high returns even with data collected from a random policy. Furthermore, with careful dynamics modeling and thorough simulation, the learned policy can more effectively handle the environmental stochasticity and generalize to states beyond the support of the offline dataset.
@@
Given a dataset, there may be various potential MDPs that behave identically on the limited set of states and actions, but their dynamics and reward functions could differ, especially on out-of-sample states and actions. This implies that we are dealing with a distribution of possible world models underlying the dataset. A common strategy in offline MBRL is to learn an ensemble of world models and treat them equally. For instance, when determining the next state, a world model is uniformly sampled from the ensemble and generate its prediction. However, different ensemble members may perform better in different regions of the state-action space, making it necessary to adapt the belief over each ensemble member based on the experience accumulated since the start of the episode. The Bayes Adaptive Markov Decision Process (BAMDP, Duff (2002)) provides a principled framework for modelling such an adaptive process. We show in Section 4.1 that, despite the need for Bayesian posterior updates, BAMDPs can still be efficiently simulated using deep ensembles. BAMCP (Guez et al. (2013)) is an efficient online planning method for solving BAMDPs. However, BAMCP has several limitations: (1) it relies on a ground-truth world model for planning; (2) it is restricted to discrete state and action spaces; and (3) its outcome is an action choice at a particular state, rather than a policy function. To address these challenges: (1) we apply a reward penalty (defined with the adapting belief) to construct a pessimistic BAMDP, preventing overexploitation of inaccurate world models (learned from the offline dataset); (2) we propose a novel planning algorithm to solve BAMDPs in continuous state and action spaces by extending BAMCP with double progressive widening (Auger et al. (2013)); and (3) we integrate the planning component as a policy improvement operator within policy iteration RL methods (Sutton & Barto (2018)), enabling the derivation of a policy suitable for real-time execution from the planning results. Specifically, the planning process is carried out through Monte Carlo Tree Search on a BAMDP. Integrating search with RL allows for the use of significantly more computation input, thereby improving policy learning performance. Grounded in the “scaling law", this paradigm has seen tremendous success in sophisticated policy learning, as demonstrated in (Silver et al. (2017b); Schrittwieser et al. (2020); AlphaProof & AlphaGeometry (2024)). Its application to offline MBRL, particularly in continuous control tasks, is a promising area for exploration.
@@
To summarize, the main contributions of this work include: (1) Introducing BAMDPs to handle model uncertainties in offline MBRL; (2) Proposing an efficient Bayes Adaptive Monte Carlo Tree Search method for planning in continuous, stochastic BAMDPs; (3) Developing the first algorithm to successfully integrate Bayesian RL, offline MBRL, and deep search for sophisticated policy learning in continuous control under highly stochastic environments; (4) Demonstrating the improvements brought by Bayesian RL and deep search across twelve D4RL MuJoCo tasks and three target tracking tasks in a stochastic tokamak control scenario (for nuclear fusion), highlighting the potential of our algorithm to tackle challenging, real-world problems.
@@ 2 Background
A Markov Decision Process (MDP, Puterman (2014)) is described as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the dynamics function and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ is the reward function, where $\Delta_{\mathcal{X}}$ denotes the set of possible probability distributions on $\mathcal{X}$. $\gamma\in[0,1)$ is a discount factor. The goal of solving an MDP is to find a policy $\pi:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$ that maximizes the expected return, defined as the cumulative discounted reward over time.
@@
A Bayes Adaptive MDP (BAMDP, Duff (2002)) can model scenarios where the precise MDP $\mathcal{M}_{\theta}=\langle\mathcal{S},\mathcal{A},\mathcal{P}_{\theta},\mathcal{R}_{\theta},\gamma\rangle$ is uncertain but is known to follow a prior distribution $b_{0}(\theta)$. During planning, a Bayes-optimal agent would update its belief over the MDP based on experience. Formally, a BAMDP can be described as a tuple $\mathcal{M}^{+}=\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\gamma\rangle$. $\mathcal{S}^{+}$ denotes the space of information states $(s,b)$, which is a composition of the physical state and the current belief over the MDP. After each transition $(s,a,r,s^{\prime})$, the belief is updated to the corresponding Bayesian posterior: $b^{\prime}(\theta)\propto b(\theta)P((s,a,r,s^{\prime})|\theta)=b(\theta)\mathcal{P}_{\theta}(s^{\prime}|s,a)\mathcal{R}_{\theta}(r|s,a)$. Accordingly, $\mathcal{P}^{+}$ and $\mathcal{R}^{+}$ can be defined as follows:
[…]
The Q-function that satisfies the Bellman optimality equations: $(\forall x=(s,b)\in\mathcal{S}^{+},a\in\mathcal{A})$
[…]
is the Bayes-optimal Q-function and $\pi^{*}(s,b)=\arg\max_{a}Q^{*}((s,b),a)$ is the Bayes-optimal policy. Actions derived from $\pi^{*}$ are executed in the real MDP and constitute the best course of action for a Bayesian agent with respect to its prior belief $b_{0}$ over the environment (Guez et al. (2014)). A BAMDP can be cast into a partially observable MDP (POMDP) [^1] Such a POMDP can be defined as a tuple $\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\mathcal{Z},\mathcal{S},\gamma\rangle$, according to its original definition in (Littman (2009)). by viewing $\mathcal{S}^{+}$ and $\mathcal{S}$ as the state and observation spaces, respectively, and setting the observation function as $\mathcal{Z}((s^{\prime},b^{\prime}),a,o)=\mathds{1}(s^{\prime}=o)$. ($\mathcal{Z}$ specifies the probability of observing $o$ under the previous action $a$ and current state $(s^{\prime},b^{\prime})$.) As a result, approaches developed for POMDPs can potentially be used to solve BAMDPs.
Bayesian Reinforcement Learning (BRL, Ghavamzadeh et al. (2015)), as introduced above, is a principled approach to dealing with uncertainty in the world model $\mathcal{M}_{\theta}$ and has two main advantages: (1) Domain knowledge can be injected by defining a proper prior belief; (2) A Bayes Adaptive policy solves the exploration-exploitation dilemma by explicitly including the belief in its state representation and incorporating belief updates into the planning process (Sorg et al. (2010)). However, Bayes-optimal planning is generally intractable, for which we introduce some approximate methods in the next section.
@@ 3 Related Works
Offline RL (Levine et al. (2020); Chen et al. (2024)) enables an agent to learn control policies from datasets of environment transitions pre-collected by a behavior policy $\mu$, i.e., $\mathcal{D}_{\mu}=\{[(s_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}]_{i=1}^{N}\}$, circumventing the need for potentially expensive or unsafe online interactions. Model-free offline RL methods directly learn value/policy functions from $\mathcal{D}_{\mu}$ but restrict the policy to remain close to $\mu$ or to behave within the support of $\mathcal{D}_{\mu}$. Offline Model-based RL (MBRL) methods, on the other hand, explicitly learn world models $\mathcal{M}_{\theta}$ from $\mathcal{D}_{\mu}$ (through supervised learning) and adopt $\mathcal{M}_{\theta}$ as a surrogate simulator, enabling the learned policy to possibly generalize to states beyond $\mathcal{D}_{\mu}$. Specifically, both planning methods (Argenson & Dulac-Arnold (2021); Zhan et al. (2022); Diehl et al. (2023)), such as MPC (Garcia et al. (1989)), and RL methods (Yu et al. (2020); Kidambi et al. (2020); Lu et al. (2022); Yu et al. (2021); Guo et al. (2022)) can be applied on top of the learned $\mathcal{M}_{\theta}$. However, since $\mathcal{D}_{\mu}$ may not span the entire state-action space, $\mathcal{M}_{\theta}$ is unlikely to be globally accurate. Learning/Planning without any safeguards against such model inaccuracy can yield poor results. In this case, the authors of (Yu et al. (2020); Kidambi et al. (2020); Lu et al. (2022)) propose learning an ensemble of world models, using ensemble-based uncertainty estimations to construct a pessimistic MDP (P-MDP), and learning a near-optimal policy atop it. Ideally, for any policy, the performance in the real environment is lower-bounded by the performance in the corresponding P-MDP (with high probability), thus avoiding being overly optimistic about an inaccurate model. For practical implementation of the P-MDP, Lu et al. (2022) conduct a thorough empirical study to compare a series of heuristics. However, none of these offline MBRL methods have modeled the problem as a BAMDP, even though Bayesian RL/planning provides a principled framework for handling model uncertainty. Additionally, with the learned model $\mathcal{M}_{\theta}$, deep and broad search in the policy learning process is possible, but the offline RL methods introduced above have not attempted such search-based policy improvement.
Monte-Carlo Tree Search (MCTS, Browne et al. (2012)) has been successfully integrated with RL, as exemplified by AlphaZero (Silver et al. (2017a)) and MuZero (Schrittwieser et al. (2020)). These methods have achieved superhuman performance in domains requiring highly precise and sophisticated decision-making processes. AlphaZero relies on given world models, whereas MuZero learns the world model and policy simultaneously by interacting with the environment. Although there have been various extensions of MuZero (Hubert et al. (2021); Schrittwieser et al. (2021); Ye et al. (2021); Danihelka et al. (2022); Antonoglou et al. (2022); Oren et al. (2022); Chen et al. (2023); Zhao et al. (2024)), most algorithms are designed for online MBRL. According to Niu et al. (2023), the applications of MuZero in offline learning, especially for continuous control in highly stochastic environments, which is our focus, still require significant improvement. Although search-based, MuZero learns a dynamics model and a reward model in a latent state space rather than the real state space, does not apply uncertainty estimation of the learned models as in usual offline MBRL methods, and does not adopt the Bayesian RL framework, making it fundamentally different from our algorithm design.
@@
As mentioned in Section 2, Bayes-optimal planning is typically intractable. Approximate methods, such as (Asmuth et al. (2009); Sorg et al. (2010); Castro & Precup (2010); Asmuth & Littman (2011); Wang et al. (2012); Fonteneau et al. (2013); Guez et al. (2013); Slade et al. (2020)) have been developed. As a representative work, BAMCP (Guez et al. (2013)) adopts MCTS for Bayes Adaptive planning and is shown to converge in probability to a near Bayes-optimal policy at the root node of the search tree. However, all these methods cannot be directly applied to large-scale MDPs with continuous state and action spaces. Guez et al. (2014) and Hoang & Vien (2020) attempt to extend the use of BAMCP to scenarios with continuous states and actions using value/policy function approximations, but they replace MCTS with simple Monte Carlo rollouts (Bertsekas & Castañón (1999)) and do not utilize UCT (Kocsis & Szepesvári (2006)). Tree search makes branches at each node in the tree, providing a more thorough look-ahead search than simple rollouts, and UCT, as an efficient exploration algorithm, is the basis for the convergence guarantee of BAMCP. Moreover, these planning algorithms are not designed for offline MBRL. How to incorporate search-based planning in the policy improvement stage of RL and how to handle the uncertainty of the learned world model require further exploration, which are shown in the following section.
@@ 4 Methodology
In this work, we propose a novel offline MBRL algorithm based on Bayes Adaptive MCTS. The core challenge is to design a Bayes Adaptive planning method that is efficient in large stochastic MDPs. In this case, we propose Continuous BAMCP in Section 4.2, which can be applied to continuous control tasks with high dynamics stochasticity. Then, in Section 4.3, we present a search-based policy iteration framework, where the search results are distilled into policy and value networks for policy improvement and policy evaluation, respectively, at each iteration. In this way, we integrate offline MBRL with Bayes Adaptive MCTS. Both components require the use of an ensemble of world models for practical implementation and uncertainty quantification, which is detailed in Section 4.1.
@@ 4.1 The Key Role of Deep Ensembles
Offline MBRL methods estimate world models $\mathcal{M}_{\theta}$ from a static dataset $\mathcal{D}_{\mu}$, which would inevitably induce epistemic uncertainty about the identity of the real MDP $\mathcal{M}^{*}$. Specifically, there could be various potential MDPs that behave identically on the limited set of states and actions in $\mathcal{D}_{\mu}$, but their dynamics and reward functions may differ, especially on out-of-sample states and actions. Thus, we are actually dealing with a distribution of world models that follow a prior distribution $b_{0}(\theta)\triangleq P(\mathcal{M}_{\theta}|\mathcal{D}_{\mu})$. As introduced in Section 2, Bayesian RL based on BAMDP is a principled framework for handling model uncertainty by explicitly including the belief over the models in its state representation. Essentially, the belief is updated with experience, providing a measure of how the models’ uncertainty has changed since the beginning of the episode. As a result, the agent can adjust its behavior upon receiving new information that reduces the epistemic uncertainty. Such an adaptive policy is necessary to act optimally in offline RL, as demonstrated in (Ghosh et al. (2022)).
The idea of deep ensembles (Lakshminarayanan et al. (2017)) is to train multiple deep neural networks as approximations of a function, each using a different weight initialization and optimized with a different mini-batch sequence. For offline MBRL, we can learn an ensemble of dynamics models $\{\mathcal{P}_{\theta}^{1},\cdots,\mathcal{P}_{\theta}^{K}\}$ and reward models $\{\mathcal{R}_{\theta}^{1},\cdots,\mathcal{R}_{\theta}^{K}\}$ [^2] In some MBRL scenarios, a certain reward function is available, for instance, as defined by domain experts. Otherwise, the reward and dynamics function $(\mathcal{R}_{\theta}^{i},\mathcal{P}_{\theta}^{i})$ are usually trained as a unified probabilistic model $\mathcal{N}(\mu_{\theta}^{i},\sigma_{\theta}^{i})$, since the reward $r$ can be viewed as an element of the next state $s^{\prime}$. from the dataset $\mathcal{D}_{\mu}$ by minimizing the following supervised learning loss: ($i=1,\cdots,K$)
$\mathcal{L}(\mathcal{P}_{\theta}^{i})=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\mu}}\left[\log\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\right],\ \mathcal{L}(\mathcal{R}_{\theta}^{i})=-\mathbb{E}_{(s,a,r)\sim\mathcal{D}_{\mu}}\left[\log\mathcal{R}_{\theta}^{i}(r|s,a)\right]$ (3)
@@
$b^{\prime}(\theta)(i)=x^{i}/\sum_{j=1}^{K}x^{j},\ x^{i}=b(\theta)(i)\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a)$ (4)
This update requires a single inference from each ensemble member, but can be parallelized for computational efficiency. Equation (4) is a practical implementation of the Bayesian posterior update based on deep ensembles, and the simplification of the definition of $b(\theta)$ makes the transitions in Bayesian RL (i.e., Equation (1)) both practical and efficient.

### v4

@@ Abstract
Offline reinforcement learning (RL) is a powerful approach for data-driven decision-making and control. Compared to model-free methods, offline model-based reinforcement learning (MBRL) explicitly learns world models from a static dataset and uses them as surrogate simulators, improving the data efficiency and enabling the learned policy to potentially generalize beyond the dataset support. However, there could be various MDPs that behave identically on the offline dataset and dealing with the uncertainty about the true MDP can be challenging. In this paper, we propose modeling offline MBRL as a Bayes Adaptive Markov Decision Process (BAMDP), which is a principled framework for addressing model uncertainty. We further propose a novel Bayes Adaptive Monte-Carlo planning algorithm capable of solving BAMDPs in continuous state and action spaces with stochastic transitions. This planning process is based on Monte Carlo Tree Search and can be integrated into offline MBRL as a policy improvement operator in policy iteration. Our “RL + Search” framework follows in the footsteps of superhuman AIs like AlphaZero, improving on current offline MBRL methods by incorporating more computation input. The proposed algorithm significantly outperforms state-of-the-art offline RL methods on twelve D4RL MuJoCo tasks and three challenging, stochastic tokamak control tasks.
@@ 1 Introduction
The success of RL typically relies on large amounts of interactions with the environment. However, in real-world scenarios, such interactions can be unsafe or costly. As an alternative, offline RL (Levine et al., 2020) leverages offline datasets of transitions, collected by a behavior policy, to train a policy. To avoid overestimation of the expected return for out-of-distribution states, which can mislead policy learning, model-free offline RL methods (Kumar et al., 2020; Wu et al., 2019) often constrain the learned policy to remain close to the behavior policy. However, acquiring a large volume of demonstrations from a high-quality behavior policy, can be expensive. This challenge has led to the development of offline model-based reinforcement learning (MBRL) approaches, such as Lu et al. (2022); Guo et al. (2022). These methods train dynamics models from offline data and optimize policies using imaginary rollouts simulated by the models. A key advantage of this approach is that the dynamics modeling process is decoupled from the behavior policy’s objectives. This makes it possible, in principle, to learn effective policies even from sub-optimal or exploratory data, so long as the dataset provides adequate coverage of the relevant state-action space.
@@
Multiple MDPs can exhibit identical behaviors on an offline dataset of states and actions, yet their underlying dynamics and reward functions may differ, particularly on out-of-sample regions. This implies that we are dealing with a distribution of possible world models underlying a dataset. A common strategy in offline MBRL is to learn an ensemble of world models and treat them equally. For instance, when determining the next state, a world model would be uniformly sampled from the ensemble to generate its prediction. However, different ensemble members may perform better in different regions of the state-action space, making it necessary to adapt the belief over each ensemble based on the experience. The Bayes Adaptive Markov Decision Process (BAMDP) (Duff, 2002) provides a principled framework for modeling such adaptations, and BAMCP (Guez et al., 2013) is an efficient online planning method for solving BAMDPs. However, BAMCP has several limitations: (1) it relies on a ground-truth world model for planning; (2) it is restricted to discrete state and action spaces; and (3) its outcome is an action choice at a particular state, rather than a policy function. To address these challenges: (1) we learn an ensemble of world models from the offline dataset and construct a pessimistic BAMDP as an estimation of the true MDP; (2) we propose a novel planning algorithm to solve BAMDPs in continuous state and action spaces based on double progressive widening (Auger et al., 2013); and (3) we distill the planning outcomes into a policy function via RL, enabling real-time execution. Notably, we provide a theoretical proof establishing the consistency of our planner in continuous, Bayes-adaptive MDP settings.
@@
Our main contributions include: (1) Firstly using BAMDPs to handle model uncertainties in offline MBRL; (2) Proposing theoretically-grounded Continuous BAMCP for planning in continuous, stochastic BAMDPs; (3) Developing the first algorithm to successfully integrate Bayesian RL, offline MBRL, and deep search for data-driven continuous control; (4) Demonstrating the state-of-the-art performance on twelve D4RL MuJoCo tasks and three target tracking tasks in tokamak control (for nuclear fusion).
@@ 2 Background
An MDP (Puterman, 2014) can be described as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the dynamics function and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ is the reward function, where $\Delta_{\mathcal{X}}$ denotes the set of probability distributions on $\mathcal{X}$. $\gamma\in[0,1)$ is a discount factor. A Bayes Adaptive MDP (BAMDP) (Duff, 2002) can model scenarios where the precise MDP $\mathcal{M}_{\theta}=\langle\mathcal{S},\mathcal{A},\mathcal{P}_{\theta},\mathcal{R}_{\theta},\gamma\rangle$ is uncertain but is known to follow a prior distribution $b_{0}(\theta)$. During planning, a Bayes-optimal agent would update its belief over the MDP based on its experience. Formally, a BAMDP can be described as a tuple $\mathcal{M}^{+}=\langle\mathcal{S}^{+},\mathcal{A},\mathcal{P}^{+},\mathcal{R}^{+},\gamma\rangle$. $\mathcal{S}^{+}$ denotes the space of information states $(s,b)$, which is a composition of the physical state and the current belief over the MDP. After each transition $(s,a,r,s^{\prime})$, the belief is updated to the corresponding Bayesian posterior: $b^{\prime}(\theta)\propto b(\theta)P((s,a,r,s^{\prime})|\theta)=b(\theta)\mathcal{P}_{\theta}(s^{\prime}|s,a)\mathcal{R}_{\theta}(r|s,a)$. Accordingly, $\mathcal{P}^{+}$ and $\mathcal{R}^{+}$ are defined as follows:
$\mathcal{P}^{+}((s^{\prime},b^{\prime\prime})|(s,b),a)=\mathds{1}(b^{\prime\prime}=b^{\prime})\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)b(\theta)d\theta$ (1)
$\mathcal{R}^{+}((s,b),a)=\int_{\theta}\mathcal{R}_{\theta}(s,a)b(\theta)d\theta$
The Q-function that satisfies the Bellman optimality equations (i.e., Eq. (2)) is the Bayes-optimal Q-function and $\pi^{*}(s,b)=\arg\max_{a}Q^{*}((s,b),a)$ is the Bayes-optimal policy. Actions derived from $\pi^{*}$ are executed in the true MDP and constitute the best course of actions for a Bayesian agent with a prior belief $b_{0}$ over the underlying MDP (Guez et al., 2014). Bayesian RL (Ghavamzadeh et al., 2015) seeks to compute the optimal policy, $\pi^{*}$, and provides a principled framework for handling uncertainty in the world model $\mathcal{M}_{\theta}$. However, the requirement for Bayesian posterior updates (i.e., Eq. (1)) often renders it computationally intractable, and we discuss several approximate methods in Section 3.
$Q^{*}(x,a)=\mathcal{R}^{+}(x,a)+\gamma\int_{x^{\prime}}V^{*}(x^{\prime})\mathcal{P}^{+}(x^{\prime}|x,a)dx^{\prime}$ (2)
$V^{*}(x^{\prime})=\max_{a}Q^{*}(x^{\prime},a),\ \forall x=(s,b)\in\mathcal{S}^{+},\ a\in\mathcal{A}$
@@ 3 Related Works
Offline model-based RL: Offline RL (Chen et al., 2024) enables an agent to learn control policies from datasets of environment transitions pre-collected by a behavior policy $\mu$, i.e., $\mathcal{D}_{\mu}=\{[(s_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}]_{i=1}^{N}\}$. Offline Model-based RL (MBRL) methods explicitly learn world models $\mathcal{M}_{\theta}$ from $\mathcal{D}_{\mu}$ and adopt $\mathcal{M}_{\theta}$ as a surrogate simulator, enabling the learned policy to possibly generalize to states beyond $\mathcal{D}_{\mu}$. Both planning methods (Argenson and Dulac-Arnold, 2021; Zhan et al., 2022; Diehl et al., 2023) and RL methods (Yu et al., 2020; Kidambi et al., 2020; Lu et al., 2022) can use the learned $\mathcal{M}_{\theta}$ to obtain a policy. However, since $\mathcal{D}_{\mu}$ may not span the entire state-action space, $\mathcal{M}_{\theta}$ is unlikely to be globally accurate. Learning/Planning without any safeguards against such model inaccuracy can yield poor results. Yu et al. (2020); Kidambi et al. (2020); Lu et al. (2022) propose learning an ensemble of world models, using ensemble-based uncertainty estimations to construct a pessimistic MDP (P-MDP), and learning a near-optimal policy atop it. Ideally, for any policy, the performance in the real environment is lower-bounded by the performance in the corresponding P-MDP (with high probability), thus avoiding being overly optimistic about an inaccurate model. Notably, none of these offline MBRL methods have modeled the problem as a BAMDP, although Bayesian RL provides a principled framework for handling model uncertainty.
@@
Bayes-adaptive planning: Approximate methods for solving BAMDPs, such as Asmuth et al. (2009); Sorg et al. (2010); Castro and Precup (2010); Asmuth and Littman (2011); Wang et al. (2012); Guez et al. (2013); Slade et al. (2020) have been developed. As a representative work, BAMCP (Guez et al., 2013) adopts Monte-Carlo Tree Search (MCTS) (Browne et al., 2012) for Bayes-adaptive planning and is shown to converge in probability to a near Bayes-optimal policy at the root node of the search tree. However, all these methods cannot be applied to large-scale MDPs with continuous state and action spaces.
Also, these planning algorithms are not designed for offline MBRL. How to incorporate planning outcomes for policy improvement in RL and how to handle the inaccuracy of the learned world model during planning still require exploration.
@@ 4 Methodology
In this section, we present a novel offline MBRL algorithm based on Bayes Adaptive MCTS. In Section 4.2, we propose a Bayes Adaptive planning method (i.e., Continuous BAMCP) that can be applied to stochastic continuous control tasks. Then, in Section 4.3, we present a search-based policy iteration framework, where the search results (from Continuous BAMCP) are distilled into policy and value networks for policy improvement/evaluation at each iteration. In this way, we integrate offline MBRL with Bayes Adaptive MCTS. Both components require the use of an ensemble of world models for either practical implementation or uncertainty quantification, as detailed in Section 4.1.
@@ 4.1 The Key Role of Deep Ensembles
Offline MBRL methods estimate world models $\mathcal{M}_{\theta}$ from a static dataset $\mathcal{D}_{\mu}$. However, there could be various MDPs that behave identically on the limited set of states and actions in $\mathcal{D}_{\mu}$, but their dynamics and reward functions may differ, especially on out-of-sample states and actions. Thus, we are actually dealing with a distribution of world models that follow a prior distribution $b_{0}(\theta)\triangleq P(\mathcal{M}_{\theta}|\mathcal{D}_{\mu})$. As introduced in Section 2, Bayesian RL handles such model uncertainty by explicitly including the belief over the models in its state representation. The belief is updated with experience, providing a measure of how the models’ uncertainty has changed since the beginning of the episode.
The idea of deep ensembles (Lakshminarayanan et al., 2017) is to train multiple deep neural networks as approximations of a function (for uncertainty quantification), each using a different weight initialization and optimized with a different mini-batch sequence. For offline MBRL, we can learn an ensemble of dynamics models $\{\mathcal{P}_{\theta}^{1},\cdots,\mathcal{P}_{\theta}^{K}\}$ and reward models $\{\mathcal{R}_{\theta}^{1},\cdots,\mathcal{R}_{\theta}^{K}\}$ from the dataset $\mathcal{D}_{\mu}$ by minimizing the following supervised learning loss: ($i=1,\cdots,K$)
$\mathcal{L}(\mathcal{P}_{\theta}^{i})=-\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\mu}}\left[\log\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\right]$ (3)
$\mathcal{L}(\mathcal{R}_{\theta}^{i})=-\mathbb{E}_{(s,a,r)\sim\mathcal{D}_{\mu}}\left[\log\mathcal{R}_{\theta}^{i}(r|s,a)\right]$
@@
$b^{\prime}(\theta)(i)\propto b(\theta)(i)\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a),\ i=1,\cdots,K$ (4)
This is a practical implementation of the Bayesian posterior (originally defined in Eq. (1)) update based on deep ensembles, where $b(\theta)$, $b^{\prime}(\theta)$, and $\mathcal{P}_{\theta}^{i}(s^{\prime}|s,a)\mathcal{R}_{\theta}^{i}(r|s,a)$ denote the prior, posterior distributions, and likelihood, respectively.